OpenStudy (anonymous):

Given: x + y = 10. If x = -21, what is y? A)11 B)-31 C)31

OpenStudy (anonymous):

I.C.P. INSANE CLOWN POSSE

OpenStudy (anonymous):

DOWN WITH THE CLOWN

OpenStudy (anonymous):

whoop whoop but i was looking for an answer lol but mwcl

OpenStudy (anonymous):

thanks

OpenStudy (triciaal):

rewrite the equation in the form y = keep the balance replace x with the given value and compute y =

OpenStudy (anonymous):

N prob bob, So you need help

OpenStudy (triciaal):

x + y = 10. y = 10 - x x = -21 y = 10 - (-21) y = 10 + 21 = 31 choice C

OpenStudy (gabylovesyou):

@triciaal is correct..

OpenStudy (anonymous):

same here it is c.

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

whoop whoop

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

Whoop whoop

OpenStudy (anonymous):

hay man see u later i going to eat lunch

OpenStudy (anonymous):

ight dueces thanks homie

OpenStudy (anonymous):

later man